Examine the function f : R→R defined by

f(x) = { (1/6) (x+1)3 x≠ 0

f(x) = { 5/6 x=0

for continuity on R. If it is not continuous at any point of R, find the nature of

discontinuity there.


1
Expert's answer
2022-01-31T17:52:48-0500

Let us examine the function "f : \R→\R" defined by


"f(x) = 
\begin{cases} \frac{1}6 (x+1)^3,\ x≠ 0\\
\frac{5}6,\ \ x=0
\end{cases}"


for continuity on "\R."

Since the elementary function "g(x)= \frac{1}6 (x+1)^3" is continuous, we conclude that the function "f" is continuous on the set "\R\setminus\{0\}." Taking into account that

"\lim\limits_{x\to 0} f(x)=\lim\limits_{x\to 0} \frac{1}6 (x+1)^3=\frac{1}6\ne\frac{5}6=f(0),"

we conclude that the function "f" has removable discontinuity at the point "x=0."